OpenStudy (anonymous):

x^2 +y^2 -3z^2 =9. Change from rectangular to spherical

OpenStudy (anonymous):

You convert using the following: \[\begin{cases}x=\rho\sin\phi\cos\theta\\ y=\rho\sin\phi\sin\theta\\ z=\rho\cos\phi\end{cases}\]